need help with this problem:4k+5-2k=3k+1, and this problem: 5-(x-3)=7x-9
Answer by sdmmadam@yahoo.com(530)

need help with this problem:4k+5-2k=3k+1, and this problem: 5-(x-3)=7x-9
1)4k+5-2k=3k+1 ----(1)
(4k-2k-3k) = 1-5 (grouping like terms,changing sign while changing side)
-k= -4
That is k=4
Verification: Putting k=4 in (1)
LHS=13 = RHS

2)5-(x-3)=7x-9 ----(2)
5-x+3 = 7x-9
(5+3+9)= 7x+x (grouping like terms,changing sign while changing side)
17 = 8x
That is 8x= 17
Therefore x = 17/8
Verification: Putting x = 17/8 in (2)
LHs = 47/8 = RHS